# View-aware Cross-modal Distillation for Multi-view Action Recognition

### v1

@@ 2 Related Work
Multi-modal Knowledge Distillation (MKD). KD [hinton2015distilling] trains a smaller or weaker student model to mimic the predictions or intermediate representations of a teacher model, improving efficiency or performance in limited conditions. Beyond matching final logits [hinton2015distilling], KD has evolved to include feature-level supervision [romero2015fitnetshintsdeepnets, zagoruyko2016paying] and relational knowledge transfer [park2019relational] to strengthen teacher-student alignment. MKD [chen2021distilling, zhang2021knowledge] extends KD to settings where the teacher and student operate over multiple modalities. The teacher typically leverages full-modality inputs, such as RGB, audio, depth, or skeleton data, while the student learns to replicate its performance under restricted modalities. Prior works have demonstrated the benefits of MKD for RGB-depth [gupta2016cross], RGB-flow [garcia2021distillation], and audio-visual [chen2021distilling, perez2020audio] tasks. In video understanding, MKD has been shown to improve recognition under missing modalities [radevski2023multimodal] by transferring temporal and cross-modal cues from the teacher to the student. In this work, we study MKD in the context of partially overlapping multi-view action recognition under missing-audio conditions. Unlike these methods, our student operates solely with visual inputs and is trained under frame-level or sequence-level supervision, the latter representing a more challenging weakly supervised setting. To address this gap, we propose the ViCoKD method with View-aware Consistency, enabling students to learn robust cross-modal and cross-view representations under the limited conditions.
3 Proposed Method
@@ 3.1 Teacher Model with Cross-modal Attention
Modality-specific Encoders. Each input sequence consists of audio data and visual data from view , which are processed by modality-specific encoders. Specifically, we use Audio Spectrogram Transformer (AST) [gong2021ast] as the audio encoder and Vision Transformer (ViT) [dosovitskiy2020image] as the visual encoder :
@@ 3.2 Teacher-to-student View-aware Cross-modal Knowledge Distillation
Feature Distillation (FD). We distill knowledge at two stages: the attended audio-visual features from the cross-modal attention, which transfer audio cues into the visual stream, and the fused representations , which integrate multi-modal context. Formally, we minimize the frame-wise Euclidean distance between the teacher’s and the student’s features as:
@@
Logit Distillation (LD). We apply logit-level distillation at the frame-level on the fused features from multiple views. Let and denote the temperature-scaled softmax outputs of the teacher and student, respectively. The loss is computed as the Kullback–Leibler divergence [kullback1951information] (KL) between these distributions as:
@@
View-aware Consistency Module. To enforce consistent predictions across different views, we introduce a view-aware consistency loss applied at the frame-level. Figure 1(b) shows the details: for each view and time step , we obtain probabilities . A binary visibility mask is generated using human detection [wang2024yolov10], where indicates that a human is visible in view at frame , suggesting a high potential for action occurrence. Given a view pair , we compute the Jensen–Shannon divergence [menendez1997jensen] (JS) between their prediction distributions:
@@ 4.1 Experimental Conditions
Data Preparation. We conduct the experiments on the publicly available MultiSensor-Home dataset [nguyen2025multisensor], which includes two home environments: Home-1 and Home-2, as shown in Figure 2. Following the official experimental setup, we use a 7:3 train/test split within each environment. To ensure a balanced class distribution, the iterative stratification strategy proposed in [sechidis2011stratification] is employed. For all experiments, we extract a fixed number of synchronized frames per sample. During training, we apply uniform sampling with slight random perturbations to generate frame sequences that cover the entire video while preserving the fixed length . This sampling strategy also augments temporal data, enhancing robustness by introducing sequence variability. For evaluation, we adopt deterministic uniform sampling without perturbation to ensure reproducibility across runs.
@@
MultiTrans [yasuda2022multi] | MultiASL [nguyen2024action] | MultiTSF [nguyen2025multisensor] | |||||
| AV V | AV V | AV V | AV V | AV V | AV V | ||||
| Home-1 | Teacher | 61.40 | 73.81 | 76.12 | |||||
| Student | 60.77 | 57.59 | 63.24 | 55.91 | 73.17 | 61.17 | |||
| LogitKD | 61.28 | 60.48 | 65.26 | 58.78 | 70.08 | 60.67 | |||
| ModalKD | 58.22 | 54.04 | 69.50 | 58.92 | 73.67 | 62.88 | |||
| DualKD | 58.84 | 55.57 | 68.42 | 60.45 | 73.12 | 63.22 | |||
| CoKD | 61.34 | 60.65 | 75.33 | 64.52 | 76.89 | 63.27 | |||
| ViCoKD (Ours) | 66.27 | 62.66 | 76.83 | 65.23 | 82.91 | 65.28 | |||
| Home-2 | Teacher | 86.60 | 90.14 | 92.12 | |||||
| Student | 84.06 | 80.83 | 82.25 | 81.67 | 86.89 | 79.16 | |||
| LogitKD | 86.70 | 85.55 | 87.59 | 84.94 | 88.60 | 84.73 | |||
| ModalKD | 85.43 | 83.35 | 91.45 | 87.98 | 90.92 | 87.56 | |||
| DualKD | 86.41 | 84.63 | 85.58 | 81.88 | 87.71 | 85.36 | |||
| CoKD | 88.64 | 86.12 | 91.77 | 86.97 | 88.42 | 87.75 | |||
| ViCoKD (Ours) | 90.86 | 89.31 | 91.80 | 89.89 | 91.27 | 89.21 | |||
@@
Models & Hyperparameters. We adopt the backbone architecture and hyperparameters of the teacher and student networks from the original work. Following Nguyen et al. [nguyen2025multisensor], we sample input videos at 2.5 FPS and fix the sequence length to frames, based on the average video duration in the dataset. The cross-modal adapter is implemented as a lightweight feed-forward module with two linear layers and a Rectified Linear Unit (ReLU) [agarap2018deep] activation in between. The human-detection masks are generated by a You Only Look Once v10 (YOLO)-based [wang2024yolov10] detector. All loss terms are equally weighted with coefficients set to for simplicity. Optimization is performed using the Adaptive moment estimation (Adam) [kingma2014adam] optimizer with an initial learning rate of , weight decay of , and a batch size of for training epochs. The learning rate is scheduled using cosine annealing [loshchilov2016sgdr] for gradual decay over the course of training.
@@ 4.2 Quantitative Results
Home-1 | Home-2 | |||
| AV V | AV V | AV V | AV V | ||
| MultiTrans | 66.27 0.59 | 62.66 0.77 | 90.86 0.36 | 89.31 0.40 | |
| MultiASL | 76.83 0.80 | 65.23 0.74 | 91.80 0.85 | 89.89 0.57 | |
| MultiTSF | 82.91 0.65 | 65.28 0.21 | 91.27 0.79 | 89.21 0.70 | |
@@
Home-1 | Home-2 | |
| MultiTrans [yasuda2022multi] | 61.40 66.31 | 86.60 88.88 | |
| MultiASL [nguyen2024action] | 73.81 75.99 | 90.14 91.78 | |
| MultiTSF [nguyen2025multisensor] | 76.12 82.16 | 92.12 92.61 |
@@ 4.3 Ablation Studies
Home-1 | Home-2 | |||
| AV V | AV V | AV V | AV V | ||
| ViCoKD | 76.83 | 65.23 | 91.80 | 89.89 | |
| View-aware Consistency Module | |||||
| w/o ConfW | 73.15 | 62.27 | 90.51 | 88.90 | |
| w/o Mask | 72.33 | 63.34 | 90.08 | 89.19 | |
| JS KL | 75.50 | 64.87 | 90.11 | 89.62 | |
| Knowledge Distillation Strategies | |||||
| w/o FD | 76.52 | 64.85 | 90.44 | 88.90 | |
| w/o LD | 76.23 | 61.51 | 92.02 | 88.76 | |
@@ 4.4 Qualitative Results
5 Limitations and Future Work
@@ Acknowledgment
Appendix A Experimental Conditions
A.1 Datasets
We conduct the experiments on the MultiSensor-Home dataset [nguyen2025multisensor], a large-scale benchmark for multi-modal and multi-view human action recognition in real-world home environments. As shown in Figure 5, each home environment is equipped with multiple synchronized RGB and audio sensors positioned to capture daily activities from different viewpoints with partial overlaps. This multi-modal multi-view setup enables both frame-level and sequence-level supervision, but also introduces challenges due to varying degrees of inter-view alignment and occlusion.
The dataset comprises two distinct home environments, denoted Home-1 and Home-2, whose statistics are summarized in Table 5. Video recordings were captured at a resolution of pixels and a frame rate of 30 FPS. The dataset consists of 5,250 untrimmed video clips, including 2,550 from Home-1 and 2,700 from Home-2, corresponding to 1,050 multi-view sequences. Each individual view contributes approximately 9.6 hours of footage in Home-1 and 11.2 hours in Home-2. With five synchronized sensors deployed in each environment, the dataset totals approximately 20.8 hours of multi-view recordings. The average recording length is 68 seconds in Home-1 and 74 seconds in Home-2. Both environments have 16 action categories (Table 6), covering a wide range of common household activities such as “enter the room”, “sit down”, “use laptop”, and “open curtain”. Some classes are environment-specific, for example, “AdjustAC” appears only in Home-1, while “WatchTV” is unique to Home-2.
| Attribute | Home-1 | Home-2 |
| Total # of Multi-view Sequences | 510 | 540 |
| Total # of Videos | 2,550 | 2,700 |
| Total Multi-view Sequence Duration | 9.6 hours | 11.2 hours |
| Total Duration | 48 hours | 56 hours |
| Average Video Duration | 68 seconds | 74 seconds |
| Average # of Actions in Video | 3 | 3 |
| # of Action Classes | 16 | 16 |
| Total # of Actions | 1,334 | 1,171 |
The two environments differ in spatial configuration and view correlation: Home-1 exhibits greater viewpoint disparity and weaker cross-view correspondence, making it more challenging for cross-view knowledge transfer, whereas Home-2 offers stronger inter-view alignment. These characteristics make MultiSensor-Home a suitable testbed for evaluating the robustness of knowledge distillation under varying degrees of view overlap and label sparsity.
| Classes | Description | #Home-1 | #Home-2 |
| AdjustAC | Adjusting air conditioner | 39 | — |
| Clean | General cleaning activity | 26 | 33 |
| CleanVacuum | Cleaning with vacuum cleaner | 48 | 45 |
| OpenCurtain | Opening the curtain | 38 | 35 |
| CloseCurtain | Closing the curtain | 39 | 35 |
| Drink | Drinking water | 51 | 59 |
| Eat | Eating food | 48 | 51 |
| Enter | Entering the room | 70 | 82 |
| Exit | Exiting the room | 88 | 82 |
| ReadBook | Reading a book | 64 | 79 |
| Sitdown | Sitting down | 247 | 142 |
| Standup | Standing up | 161 | 126 |
| TurnOnLamp | Turning on the lamp | 57 | 49 |
| TurnOffLamp | Turning off the lamp | 52 | 41 |
| UseLaptop | Using a laptop computer | 196 | 114 |
| UsePhone | Using a phone | 110 | 102 |
| WatchTV | Watching the television | — | 96 |
| Total | 1,334 | 1,171 | |
A.2 Evaluation Metrics
Following [kobayashi2023two, nguyen2024action, nguyen2025multisensor], we evaluate performance using the mean Average Precision (mAP) based on a macro-averaged metric, which is widely used in multi-label classification tasks. mAP is computed independently for each class over samples and then aggregated across all classes as:
where AP denotes the Average Precision computed for a set of predictions and ground truths for class . This metric effectively evaluates how well the model ranks relevant instances ahead of irrelevant ones across multiple labels.
A.3 Resource
All experiments are conducted on a FUJITSU CX2570M5 server equipped with two Intel Xeon Gold 6230 CPUs and four NVIDIA Tesla V100 GPUs.
Appendix B Experimental Results
In this section, we provide additional experimental results to further evaluate the effectiveness of the proposed Cross-modal Knowledge Distillation (ViCoKD) method.
B.1 Testing mAP Curves under Different Distillation Settings
Figures 6 and 7 present the testing mAP curves for the MultiTrans [yasuda2022multi] and MultiTSF [nguyen2025multisensor] backbones under various distillation strategies. A consistent trend emerges across both architectures: the proposed ViCoKD method converged more rapidly and attained higher or more stable mAP compared to LogitKD and CoKD. In particular, under the AV V setting, ViCoKD demonstrated clear advantages by surpassing the non-distilled student early in training and sustaining strong performance throughout.
For MultiTrans [yasuda2022multi] backbone (Figure 6), the improvements were most noticeable in the Home-1 environment, where other methods showed slower convergence or larger fluctuations. For MultiTSF [nguyen2025multisensor] backbone (Figure 7), ViCoKD not only delivered the best final performance but also reduced variance across epochs, suggesting more stable optimization dynamics.
Overall, these additional results confirm that the effectiveness of ViCoKD is not limited to a single backbone but generalized well across different temporal fusion architectures. This reinforces the role of View-aware Consistency in guiding robust representation learning under partial cross-modal supervision.
B.2 Qualitative Results
In addition to the qualitative results presented in the main text, Figure 8 provides further visualizations of attention maps for diverse actions. Across both Home-1 and Home-2 environments, the proposed ViCoKD method consistently produced human and the action-related attention than either the baseline student or the teacher.
Home-1 environment: For the “sit down” action (Figure 8(a)) across Views 1, 2, and 3, the teacher model distributed attention over the actor’s body posture changes, while the baseline student often drifted toward background elements such as furniture or walls. ViCoKD, in contrast, narrowed its focus directly on the seated human across all views, maintaining cross-view consistency. Similarly, in the “clean by vacuum” action (Figure 8(c)) on Views 4 and 5, the teacher highlighted the person in View 4 and the vacuum cleaner in View 5. ViCoKD recovered this alignment by highlighting both the human and the vacuum cleaner in both views, reliably centering on the human and the action-related tool across perspectives.
Home-2 environment: For the “watch TV” action (Figure 8(b)), the teacher mainly attended to the TV screen but often neglected the human actor, especially in View 5. The baseline student, in contrast, produced scattered and weak attention maps, frequently drifting toward irrelevant floor regions. ViCoKD recovered a more balanced focus, consistently attending to both the TV and the human subject across all views, thereby capturing the essential human-object interaction that defines the action. For the “use laptop” action (Figure 8(d)), the attention maps of the teacher, baseline student, and ViCoKD appear more similar compared to other actions. All three models highlighted the subject and the laptop region to some degree, reflecting the strong visual saliency of the interaction.

### v2

@@ 2 Related Work
Multi-modal Knowledge Distillation (MKD). KD [hinton2015distilling] trains a smaller or weaker student model to mimic the predictions or intermediate representations of a teacher model, improving efficiency or performance in limited conditions. Beyond matching final logits [hinton2015distilling], KD has evolved to include feature-level supervision [romero2015fitnetshintsdeepnets, zagoruyko2016paying] and relational knowledge transfer [park2019relational] to strengthen teacher–student alignment. MKD [chen2021distilling, zhang2021knowledge] extends KD to settings where the teacher and student operate over multiple modalities. The teacher typically leverages full-modality inputs, such as RGB, audio, depth, or skeleton data, while the student learns to replicate its performance under restricted modalities. Prior works have demonstrated the benefits of MKD for RGB-depth [gupta2016cross], RGB-flow [garcia2021distillation], and audio-visual [chen2021distilling, perez2020audio] tasks. In video understanding, MKD has been shown to improve recognition under missing modalities [radevski2023multimodal] by transferring temporal and cross-modal cues from the teacher to the student. In this work, we study MKD in the context of partially overlapping multi-view action recognition under missing-audio conditions. Unlike these methods, our student operates solely with visual inputs and is trained under frame-level or sequence-level supervision, the latter representing a more challenging weakly supervised setting. To address this gap, we propose the ViCoKD method with View-aware Consistency, enabling students to learn robust cross-modal and cross-view representations under the limited conditions.
3 ViCoKD: View-aware Cross-modal Distillation Framework
@@ 3.1 Teacher Model with Cross-modal Attention
Modality-specific Encoders. Each input sequence consists of audio data and visual data from view , which are processed by modality-specific encoders. Specifically, we use Audio Spectrogram Transformer (AST) [gong2021ast] as the audio encoder and Vision Transformer (ViT) [dosovitskiy2020image] as the visual encoder as:
@@ 3.2 Teacher-to-student View-aware Cross-modal Knowledge Distillation
Feature Distillation (FD). We distill knowledge at two stages: (1) The attended audio-visual features from the cross-modal attention, which transfer audio cues into the visual stream, and (2) The fused representations , which integrate multi-modal context. Formally, we minimize the frame-wise Euclidean distance between the teacher’s and the student’s features as:
@@
Logit Distillation (LD). We apply logit-level distillation at the frame-level on the fused features from multiple views. Let and denote the temperature-scaled softmax outputs of the teacher and student, respectively. The loss is computed as the Kullback–Leibler (KL) divergence [kullback1951information] between these distributions as:
@@
View-aware Consistency Module. To enforce consistent predictions across views, we introduce a view-aware consistency loss applied at the frame-level. Figure 1(b) shows the details: For each view at time step , we obtain probabilities . A binary visibility mask is generated using human detection [wang2024yolov10], where indicates that a human is visible in view at frame , suggesting a high potential for action occurrence. Given a view pair , we compute the Jensen–Shannon (JS) divergence [menendez1997jensen] between their prediction distributions as:
@@ 4.1 Experimental Conditions
Data Preparation. We conduct the experiments on the publicly available MultiSensor-Home dataset [nguyen2025multisensor], which includes two home environments: Home-1 and Home-2, as shown in Figure 2. Following the official experimental setup, we use a 7:3 train/test split within each environment. To ensure a balanced class distribution, the iterative stratification strategy proposed by Sechidis et al. [sechidis2011stratification] is employed. For all experiments, we extract a fixed number of synchronized frames per sample. During training, we apply uniform sampling with slight random perturbations to generate frame sequences that cover the entire video while preserving the fixed length . This sampling strategy also augments temporal data, enhancing robustness by introducing sequence variability. For evaluation, we adopt deterministic uniform sampling without perturbation to ensure reproducibility across runs.
@@
| Environment | Method | MultiTrans [yasuda2022multi] | MultiASL [nguyen2024action] | MultiTSF [nguyen2025multisensor] | |||
| AV V | AV V | AV V | AV V | AV V | AV V | ||
| Home-1 | Teacher | 61.40 | 73.81 | 76.12 | |||
| Student | 60.77 | 57.59 | 63.24 | 55.91 | 73.17 | 61.17 | |
| LogitKD | 61.28 | 60.48 | 65.26 | 58.78 | 70.08 | 60.67 | |
| ModalKD | 58.22 | 54.04 | 69.50 | 58.92 | 73.67 | 62.88 | |
| DualKD | 58.84 | 55.57 | 68.42 | 60.45 | 73.12 | 63.22 | |
| CoKD | 61.34 | 60.65 | 75.33 | 64.52 | 76.89 | 63.27 | |
| ViCoKD (Ours) | 66.27 | 62.66 | 76.83 | 65.23 | 82.91 | 65.28 | |
| Home-2 | Teacher | 86.60 | 90.14 | 92.12 | |||
| Student | 84.06 | 80.83 | 82.25 | 81.67 | 86.89 | 79.16 | |
| LogitKD | 86.70 | 85.55 | 87.59 | 84.94 | 88.60 | 84.73 | |
| ModalKD | 85.43 | 83.35 | 91.45 | 87.98 | 90.92 | 87.56 | |
| DualKD | 86.41 | 84.63 | 85.58 | 81.88 | 87.71 | 85.36 | |
| CoKD | 88.64 | 86.12 | 91.77 | 86.97 | 88.42 | 87.75 | |
| ViCoKD (Ours) | 90.86 | 89.31 | 91.80 | 89.89 | 91.27 | 89.21 | |
@@
Models & Hyperparameters. We adopt the backbone architecture and hyperparameters of the teacher and student networks from the original work. Following Nguyen et al. [nguyen2025multisensor], we sample input videos at 2.5 FPS and fix the sequence length to frames, based on the average video duration in the dataset. The cross-modal adapter is implemented as a lightweight feed-forward module with two linear layers and a Rectified Linear Unit (ReLU) [agarap2018deep] activation in between. The human-detection masks are generated by a You Only Look Once (YOLO) v10-based [wang2024yolov10] detector. All loss terms are equally weighted with coefficients set to for simplicity. Optimization is performed using the Adaptive moment estimation (Adam) [kingma2014adam] optimizer with an initial learning rate of , weight decay of , and a batch size of for training epochs. The learning rate is scheduled using cosine annealing [loshchilov2016sgdr] for gradual decay over the course of training.
@@ 4.2 Quantitative Results
| Backbone | Home-1 | Home-2 | ||
| AV V | AV V | AV V | AV V | |
| MultiTrans | 66.27 0.59 | 62.66 0.77 | 90.86 0.36 | 89.31 0.40 |
| MultiASL | 76.83 0.80 | 65.23 0.74 | 91.80 0.85 | 89.89 0.57 |
| MultiTSF | 82.91 0.65 | 65.28 0.21 | 91.27 0.79 | 89.21 0.70 |
@@
| Backbone | Home-1 | Home-2 |
| MultiTrans [yasuda2022multi] | 61.40 66.31 | 86.60 88.88 |
| MultiASL [nguyen2024action] | 73.81 75.99 | 90.14 91.78 |
| MultiTSF [nguyen2025multisensor] | 76.12 78.25 | 92.12 92.61 |
@@ 4.3 Ablation Studies
| Method | Home-1 | Home-2 | ||
| AV V | AV V | AV V | AV V | |
| ViCoKD | 76.83 | 65.23 | 91.80 | 89.89 |
| View-aware Consistency Module | ||||
| w/o ConfW | 73.15 | 62.27 | 90.51 | 88.90 |
| w/o Mask | 72.33 | 63.34 | 90.08 | 89.19 |
| JS KL | 75.50 | 64.87 | 90.11 | 89.62 |
| Knowledge Distillation Strategies | ||||
| w/o FD | 76.52 | 64.85 | 90.44 | 88.90 |
| w/o LD | 76.23 | 61.51 | 92.02 | 88.76 |
@@ 4.4 Qualitative Results
4.5 Limitations and Future Work